(a)The table below show Business statistics I examination marks for a class of one hundred students in Mount Kenya university.

Marks

10-19

20-29

30-39

40-49

50-59

60-69

Number of students

4

8

a

22

48

b

The mean score of the students was 46.5

         (i) Show that the values of a and b are 12 and 6 respectively.                       (5marks)

         (ii) Estimate the median for the above sample.                                               (3marks)

        (iii) Determine the variance and the standard deviation                                  (4 marks)     

        (iv) Given that 32% of the students failed the examination, what was the cut off points.

(3 marks)

(v)Estimate the mode for this data.                                                                (3marks)



1
Expert's answer
2021-10-24T17:33:27-0400

i)

mean:

"\mu=\dfrac{\sum x_i}{n}=\dfrac{14.5\cdot4+24.5\cdot8+34.5a+44.5\cdot22+54.5\cdot48+64.5b}{4+8+a+22+48+b}=46.5"


"n=4+8+a+22+48+b=100"


Then,

"a+b=18"

"\implies3849+34.5a+64.5b=4650"

"\implies34.5(18-b)+64.5b=801"

"\implies30b=180"

"b=180/30=6,\ a=18-6=12"

"\boxed{a=12}\ \ \boxed{b=6}"



ii)

"median=l + [(n\/2\u2212c)\/f] \u00d7 h"

where,

  • l = lower limit of median class
  • n = total number of observations
  • c = cumulative frequency of the preceding class
  • f = frequency of median class
  • h = class size


cumulative frequency:

10-19: 4

20-29: 12

30-39: 24

40-49: 46

50-59: 94

60-69: 100


median class: 50-59


"median=50+[(100/2-46)/94]\cdot9=50.38"


iii)

variance:

"\sigma^2=\dfrac{\sum f_iM_i^2-n\mu^2}{n-1}"

where fi is frequency of i-th range,

Mi is midpoint of i-th range.


"\sigma^2=\dfrac{4\cdot14.5^2+8\cdot24.5^2+12\cdot34.5^2+22\cdot44.5^2+48\cdot54.5^2+6\cdot64.5^2-100\cdot46.5^2}{99}\\\ \\=\dfrac{231025-216225}{99}=149.49"


standard deviation:


"\sigma=\sqrt{149.49}=12.23"


iv)

32% percentile:

"P_{32}=l+w(32-c)\/f"

where l is the lower boundary of the class containing P32,

w is the class interval size of the class containing P32,

f is the frequency of the class containing P32,

c is the cumulative frequency of the class preceding the class containing P32


class containing P32: 40-49

Then:

"P_{32}=40+9(32-24)\/22=43.27"


v)

"mode=l+\dfrac{f_m-f_1}{(f_m-f_1)(f_m-f_)}\cdot h"

where

l is the lower limit of the modal class,

fm is the frequency of the modal class,

f1 is the frequency of class preceding the modal class,

f2 is the frequency of class succeeding the modal class,

h is the width of the modal class.


modal class (a class that has the highest frequency): 50-59


"mode=50+\dfrac{48-22}{48-22+48-6}\cdot9=53.44"